# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

### v2

@@ Appendix C Gradient and Hessian matrix Derivation of $\widetilde{f}_{1}$
$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{y,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60b)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60c)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}\partial p_{x,n}}$
$=\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60d)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right).$ (60e)
@@
By substituting the results of the first-order and second-order derivatives to (39) and (40), the corresponding gradient and the Hessian matrix of $\widetilde{f}_{1}$ will be obtained.
This finishes the derivations.

### v3

@@ Appendix C Gradient and Hessian matrix Derivation of $\widetilde{f}_{1}$
$\!\!\!\!\!\begin{bmatrix}\!\!\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}^{2}}&\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}\partial p_{y,n}}\\ \frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}\partial p_{x,n}}&\!\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}^{2}}\end{bmatrix}\!\!=\!\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\mathbf{K}_{k}\cos\!\left(\overline{\tau}_{n,k}\right)\!,$ (54b)
where $\overline{\tau}_{n,k}=\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}$.
@@
By substituting the above results to (36) and (37), the corresponding gradient and the Hessian matrix of $\widetilde{f}_{1}$ will be obtained. This finishes the derivations.